# Adversarial Prompt Tuning for Vision-Language Models

### v2

@@ Abstract
With the rapid advancement of multimodal learning, pre-trained Vision-Language Models (VLMs) such as CLIP have demonstrated remarkable capacities in bridging the gap between visual and language modalities. However, these models remain vulnerable to adversarial attacks, particularly in the image modality, presenting considerable security risks. This paper introduces Adversarial Prompt Tuning (AdvPT), a novel technique to enhance the adversarial robustness of image encoders in VLMs. AdvPT innovatively leverages learnable text prompts and aligns them with adversarial image embeddings, to address the vulnerabilities inherent in VLMs without the need for extensive parameter training or modification of the model architecture. We demonstrate that AdvPT improves resistance against white-box and black-box adversarial attacks and exhibits a synergistic effect when combined with existing image-processing-based defense techniques, further boosting defensive capabilities. Comprehensive experimental analyses provide insights into adversarial prompt tuning, a novel paradigm devoted to improving resistance to adversarial images through textual input modifications, paving the way for future robust multimodal learning research. These findings open up new possibilities for enhancing the security of VLMs. Our code is available at https://github.com/jiamingzhang94/Adversarial-Prompt-Tuning.
@@ 1 Introduction
In the image domain, adversarial training has been proven to be the most effective approach for training robust DNNs against adversarial examples (inputs with adversarial noise) [18]. Adversarial training is usually formulated as a min-max optimization problem, which generates adversarial examples at each training iteration to update the image encoder. Therefore, it is computationally expensive, and thus cannot be easily applied to train large VLMs. As such, recent works turn to input pre-processing techniques like diffusion-based purification to improve the adversarial robustness of VLMs [37, 19, 20].
@@
We evaluate AdvPT against both white-box and black-box adversarial attacks on 8 image datasets, and show that it outperforms the vanilla CLIP (with hand-crafted prompts) by a considerable margin. By focusing on textual input processing and alignment, AdvPT opens up a new direction for augmenting adversarial robustness in VLMs. It can also be integrated with image-based defense strategies to further boost the adversarial robustness of the image modality. We also observe a generalization-robustness trade-off in AdvPT, similar to that in traditional adversarial training. We further evaluate the domain transferability of the learnable vectors, testing their performance across various datasets after training on a specific one. Lastly, we conduct an in-depth analysis of the learned vectors and reveal the closest words associated with the vectors, gaining more understanding of the working mechanism of AdvPT.
@@ 2.3 Adversarial Defenses
Combating against adversarial images remains an unresolved challenge. Adversarial defenses broadly fall into two camps: model robustification methods and input denoising methods. The former includes methods like adversarial training [18], TRADES [40] and MART [34]. This methodology is usually expressed as a min-max optimization problem, with continuous updates to the model parameters across all training iterations. However, this process is computationally demanding, posing difficulties for deployment on VLMs due to the scale of the model and dataset. As a result, the latter approach based on the image process has emerged as a solution suited for VLMs.
The adversarial defense through input image modification is straightforward in its essence. It removes or weakens the impact of adversarial noise through inference-time methods such as input transformations [7, 19], smoothing [16], and rescaling [37]. For example, Xie et al. [37] employed random image rescaling to diminish adversarial effects, and Mustafa et al. [19] utilized image super-resolution as a defense mechanism. Although somewhat limited in efficacy, these methods are pragmatically valuable for their efficiency. Recently, adversarial purification based on diffusion models has emerged [39, 20]. Nie et al.[20] introduced the powerful adversarial purification, DiffPure, to address the shortcomings of previous approaches, albeit with increased time complexity.
@@ 3 Revisiting Clip and the Adversarial Robustness of Its Image Encoder
Adversarial Robustness of CLIP’s Image Encoder
We first introduce our threat model, which describes the assumed knowledge of the adversary, from what inputs they can manipulate to their access to the model architecture and parameters. Our study focuses on the adversarial robustness of image encoders, assuming that the attacker has full knowledge of the model architecture and parameters, and can perturb the image input. However, the adversary has no control over the textual input. Therefore, both white-box and black-box adversarial attacks targeting the image encoder can be used as pertinent evaluative tools in this context. Nevertheless, potent adaptive attacks like AutoAttack are outside of our research scope [3, 30], as they rely on a classification layer to craft adversarial examples. Similarly, multimodal adversarial attacks integrated with text adversarial attacks are also not applicable here [42, 45, 41].
We now introduce the adversarial attacks that target the image encoders. Consider an original input image $x$, with $\delta$ symbolizing adversarial noise. The adversarial example $x^{\prime}=x+\delta$, once processed by the image encoder $E$, generates an adversarial embedding $e^{\prime}$. The adversarial embedding $e^{\prime}$ markedly diverges from the embedding $e$ of the original image, impairing the accuracy of matching with textual descriptions. In this work, we focus on $\ell_{\infty}$-norm constrained perturbations, where each $\delta$ adheres to $\|\delta\|_{\infty}\leq\epsilon$, with $\epsilon$ denoting the maximum allowable perturbation magnitude.
To defend against adversarial images, existing defense methods generally fall into two categories: model robustification methods and input denoising methods. As mentioned above, model robustification methods like adversarial training struggle to handle VLMs due to efficiency issues, rendering input denoising methods like diffusion-based denoising a practical approach to strength VLMs. The input denoising operation can be conceptualized as a function $h$ that processes adversarial images, aiming to minimize the disparity between $E(h(x^{\prime}))$ and $E(x)$.
@@ 4 Adversarial Prompt Tuning
Overview
Our proposed method, AdvPT, involves optimizing learnable vectors as text prompts to enhance the robustness against image adversarial attacks. This diverges from previous context optimization approaches [44, 43] aimed at increasing image recognition rates. Fig. 2 provides an framework overview of AdvPT. On a $K$-class dataset $D=\{(x_{i},y_{i})\}_{i=1}^{N}$ of $N$ images and corresponding texts, AdvPT begins with feeding the clean images $x$ into the image encoder $E$ to generate its adversarial image $x^{\prime}$. The adversarial images are then fed into the image encoder $E$ to obtain the adversarial image embeddings into an adversarial embedding bank $\mathbf{A}\in\mathbb{R}^{N\times L}$, where $L$ is the embedding dimension (512 for CLIP). The image encoder $E$ is discarded in the subsequent steps. In other words, the adversarial embeddings are not generated in real-time during training, which is in sharp contrast to traditional adversarial training [18].
@@ 4.1 Adversarial Embedding Bank Generation
To improve the image encoder $E$’s adversarial robustness, AdvPT first generates adversarial images on encoder $E$, then re-feeds them into the encoder to obtain and store their adversarial embeddings. Note that AdvPT differs greatly from adversarial training, which iteratively generates adversarial examples at each iteration of training and continuously updates the target model on the generated adversarial examples, leading to significant computational costs. Conversely, AdvPT fixes the parameters of the image encoder $E$, channeling focus exclusively on updating the learnable vectors at the input of the text encoder $G$. This strategy significantly diminishes the number of learnable parameters. With the image encoder $E$ frozen, the generation of the adversarial examples is only a one-pass process. These examples, once processed through $E$, constitute the adversarial embedding bank $\mathbf{A}$. After this step, the image encoder $E$ is discarded, leaving only the adversarial embedding bank $\mathbf{A}$ for the subsequent prompt tuning.
@@
The design of the adversarial embedding bank presents significant advantages. Primarily, it eliminates the need for redundant forward and backward passes through the image encoder, thereby greatly saving computational time. Moreover, the embedding space’s lower dimensionality (512) compared to the original image space (224$\times$224) substantially reduces the required computational memory.
@@ 4.2 Learnable Vector Optimization
The next phase in AdvPT involves the construction and optimization of the learnable vectors. Specifically, our method seeks to model textual prompts with learnable vectors $V=[v_{1},v_{2},\ldots,v_{M}]$, optimized by aligning them with adversarial embeddings, thus rectifying the non-robust features of the images utilized by the model. Initially, the text prompts $[v_{1},v_{2},\ldots,v_{M},c_{i}]_{i=1}^{K}$ are fed into the text encoder $G$, producing text embeddings $\mathbf{T}=[w_{1},w_{2},\dots,w_{K}]^{T}\in\mathbb{R}^{L\times K}$. In the fine-tuning phase, each iteration retrieves a mini-batch $\mathbf{B}=[e^{\prime}_{1},e^{\prime}_{2},\dots,e^{\prime}_{b}]\in\mathbb{R}^{b\times L}$ from the adversarial embeddings bank $\mathbf{A}$. Subsequently, the similarity scores $\mathbf{S}=\mathbf{B}\mathbf{T}\in\mathbb{R}^{b\times K}$ can be calculated, with each element representing the prediction score in the following manner:
@@
Table 1: Accuracy (%) under white-box PGD-40 and black-box RAP attacks: The “( ↑)” indicates the margin by which AdvPT surpasses the vanilla CLIP, i.e., CLIP using hand-crafted prompts.
@@
Table 2: Accuracy (%) under white-box PGD-40 Attack: The “+AdvPT” indicates our method combined with the input denoising defense, where “( ↑)” denotes improved performance. The best results are highlighted in bold.
@@
5 Experiments
In this section, we begin by comparing the adversarial robustness of our proposed approach with hand-crafted prompts under both white-box and black-box adversarial attacks. Second, we compare our method with the state-of-the-art input denoising defensive approaches. Additionally, we investigate the trade-off between generalizability and adversarial robustness in the context of prompt tuning. We also discuss the efficiency between our method and the conventional adversarial training. Next, we examine the performance of learnable vectors when trained on a specific dataset but evaluated across various distinct datasets. Finally, we carry out an experimental analysis into interpreting the learnable vectors and perform an exhaustive analysis of hyperparameters.
5.1 Experimental Settings
Datasets
We conduct our study mainly on 8 high-resolution vision datasets: Pets [23], Flowers [21], ImageNet [27], Food101 [1], SUN397 [36], DTD [2], EuroSAT [9], and UCF101 [28]. We adhered to the division of training and testing sets as established in the setup of [44]. For the ImageNet test set, in a manner consistent with prior studies focusing on adversarial attacks [5, 33, 38], we use 1,000 images which are randomly sampled (one image per class). Furthermore, to assess the domain generalization capabilities of learnable vectors, we employed ImageNet as the source dataset and selected four variant datasets, namely ImageNetV2 [26], ImageNet-Sketch [32] (sampled 1,000 images in the same manner as ImageNet), ImageNet-A [11], and ImageNet-R [10], each presenting unseen domain shifts, as the target datasets for evaluation.
Models
Adversarial Attacks
To evaluate adversarial robustness for image encoder, we introduced both white-box and black-box adversarial attacks. For white-box adversarial attacks, we employed PGD-40 as the state-of-the-art method suitable for this context. Regarding black-box attacks, we implemented state-of-the-art black-box attack RAP [24].
Adversarial Defenses
Implementation Details
Our methodology builds upon the CoOp framework [^1] https://github.com/KaiyangZhou/CoOp. Our training process consists of 5 epochs with a batch size of 512 on ImageNet, and 100 epochs with a batch size of 32 on other datasets. The learnable vectors are optimized via SGD, starting with an initial learning rate of 0.002 for ViT-L/14 and 0.005 for ViT-B/16, and adjusted by cosine annealing. The number of learnable vector $M=32$. To construct the adversarial embedding bank $\mathbf{A}$, we apply the PGD-10 attack with a maximum perturbation of 8/255 over 10 iterations. For white-box adversarial attack on the test set, we utilize PGD-40 with a maximum perturbation of 16/255 over 40 iterations. We conduct black-box adversarial attacks on the test set using RAP for 400 iterations. For the selection of surrogate model of RAP attack, we employ ResNet-50 with torchvision weights for ImageNet, and adopt CLIP’s ResNet-50 as the backbone with training an additional fully connected layer for other datasets. The hyperparameter $\sigma$ in Super-resolution was set to 0.2. The pre-trained diffusion models in DiffPure is Guided Diffusion [4] and the time step was set to 150.
Table 3: AdvPT vs. vanilla CLIP on distribution shift.

### v3

@@ Abstract
With the rapid advancement of multimodal learning, pre-trained Vision-Language Models (VLMs) such as CLIP have demonstrated remarkable capacities in bridging the gap between visual and language modalities. However, these models remain vulnerable to adversarial attacks, particularly in the image modality, presenting considerable security risks. This paper introduces Adversarial Prompt Tuning (AdvPT), a novel technique to enhance the adversarial robustness of image encoders in VLMs. AdvPT innovatively leverages learnable text prompts and aligns them with adversarial image embeddings, to address the vulnerabilities inherent in VLMs without the need for extensive parameter training or modification of the model architecture. We demonstrate that AdvPT improves resistance against white-box and black-box adversarial attacks and exhibits a synergistic effect when combined with existing input denoising defense techniques, further boosting defensive capabilities. Comprehensive experimental analyses provide insights into adversarial prompt tuning, a novel paradigm devoted to improving resistance to adversarial images through textual input modifications, paving the way for future robust multimodal learning research. These findings open up new possibilities for enhancing the security of VLMs. Our code is available at https://github.com/jiamingzhang94/Adversarial-Prompt-Tuning. footnotetext: Corresponding authors: Xingjun Ma and Jitao Sang.
Keywords:
Adversarial defense Vision-Language models Prompt tuning
@@ 1 Introduction
In the image domain, adversarial training (AT) has been proven to be the most effective approach for training robust DNNs against adversarial examples (inputs with adversarial noise) [18]. AT is usually formulated as a min-max optimization problem, which generates adversarial examples at each training iteration to update the image encoder. Therefore, it is computationally expensive, and thus cannot be easily applied to train large VLMs. As such, recent works turn to input pre-processing techniques like diffusion-based purification to improve the adversarial robustness of VLMs [39, 20, 21].
@@
We evaluate AdvPT against both white-box and black-box adversarial attacks on 8 image datasets, and show that it outperforms the vanilla CLIP (with hand-crafted prompts) by a considerable margin. By focusing on textual input processing and alignment, AdvPT opens up a new direction for augmenting adversarial robustness in VLMs. It can also be integrated with image-based defense strategies to further boost the adversarial robustness of the image modality. We also observe a generalization-robustness trade-off in AdvPT, similar to that in traditional AT. We further evaluate the domain transferability of the learnable vectors, testing their performance across various datasets after training on a specific one. Lastly, we conduct an in-depth analysis of the learned vectors and reveal the closest words associated with the vectors, gaining more understanding of the working mechanism of AdvPT.
@@ 2.3 Adversarial Defenses
Combating against adversarial images remains an unresolved challenge. Adversarial defenses broadly fall into two camps: model robustification methods and input denoising methods. The former includes methods like AT [18], Fast AT [37], TRADES [42] and MART [35]. This methodology is usually expressed as a min-max optimization problem, with continuous updates to the model parameters across all training iterations. However, this process is computationally demanding, posing difficulties for deployment on VLMs due to the scale of the model and dataset. As a result, the latter approach based on the image process has emerged as a solution suited for VLMs.
The adversarial defense through input image modification is straightforward in its essence. It removes or weakens the impact of adversarial noise through inference-time methods such as input transformations [7, 20], smoothing [16, 29, 44], and rescaling [39]. For example, Xie et al. [39] employed random image rescaling to diminish adversarial effects, and Mustafa et al. [20] utilized image super-resolution as a defense mechanism. Although somewhat limited in efficacy, these methods are pragmatically valuable for their efficiency. Recently, adversarial purification based on diffusion models has emerged [41, 21]. Nie et al.[21] introduced the powerful adversarial purification, DiffPure, to address the shortcomings of previous approaches, albeit with increased time complexity. Mao et al.[19] identifies that AT of the CLIP on one dataset struggles to impact another dataset, defining this as the zero-shot adversarial robustness problem, and introduced visual prompt tuning [13] to address this.
@@ 3.1 CLIP
3.2 Adversarial Robustness of CLIP’s Image Encoder
We first introduce our threat model, which describes the assumed knowledge of the adversary, from what inputs they can manipulate to their access to the model architecture and parameters. Our study focuses on the adversarial robustness of image encoders, assuming that the attacker has full knowledge of the model architecture and parameters of image and text encoders, and can perturb the image input. However, the adversary has no control over the textual input nor knowledge of prompt tuning. Therefore, text adversarial attacks are also not applicable here [45, 48, 43].
We now introduce the adversarial attacks that target the image encoders. Consider an original input image $x$, with $\delta$ symbolizing adversarial noise. The adversarial example $x^{\prime}=x+\delta$, once processed by the image encoder $E$, generates an adversarial embedding $e^{\prime}$.
Adversaries can employ two objective functions to impair the accuracy of matching with textual descriptions. The first objective is to make the adversarial embedding $e^{\prime}$ markedly diverge from the embedding $e$ of the original image, i.e., to maximize the discrepancy between $e^{\prime}$ and $e$. The second objective is to ensure the adversarial embedding $e^{\prime}$ does not align with the corresponding ground-truth textual description embedding $w_{g}$, i.e., to maximize the discrepancy between $e^{\prime}$ and $w_{g}$. PGD and AutoAttack are deployed to represent the former and latter objectives, respectively. In this work, we focus on $\ell_{\infty}$-norm constrained perturbations, where each $\delta$ adheres to $\|\delta\|_{\infty}\leq\epsilon$, with $\epsilon$ denoting the maximum allowable perturbation magnitude.
To defend against adversarial images, existing defense methods generally fall into two categories: model robustification methods and input denoising methods. As mentioned above, model robustification methods like AT struggle to handle VLMs due to efficiency issues. The input denoising operation can be conceptualized as a function $h$ that processes adversarial images, aiming to minimize the disparity between $E(h(x^{\prime}))$ and $E(x)$.
@@ 4 Adversarial Prompt Tuning
Figure 2: An overview of the AdvPT framework.
Overview.
Our proposed method, AdvPT, involves optimizing learnable vectors as text prompts to enhance the robustness against image adversarial attacks. This diverges from previous context optimization approaches [47, 46] aimed at increasing image recognition rates. Fig. 2 provides an framework overview of AdvPT. On a $K$-class dataset $D=\{(x_{i},y_{i})\}_{i=1}^{N}$ of $N$ images and corresponding texts, AdvPT begins with feeding the clean images $x$ into the image encoder $E$ to generate its adversarial image $x^{\prime}$. The adversarial images are then fed into the image encoder $E$ to obtain the adversarial image embeddings into an adversarial embedding bank $\mathbf{A}\in\mathbb{R}^{N\times L}$, where $L$ is the embedding dimension. The image encoder $E$ is discarded in the subsequent steps. This approach is entirely distinct from traditional defensive methods, which, whether through augmentation (e.g., visual prompt tuning [13, 19]) or modification (e.g., AT [18]) of the parameters of CLIP’s image encoder branch, rely on on-the-fly adversarial example generation during each training epoch. Even with partial parameter tuning (visual prompt), the adversarial example generation necessitates complete forward and backward propagation of gradients through the image encoder, resulting in an untenable burden in the context of VLMs.
@@ 4.1 Adversarial Embedding Bank Generation
To improve the image encoder $E$’s adversarial robustness, AdvPT first generates adversarial images on encoder $E$, then re-feeds them into the encoder to obtain and store their adversarial embeddings. Note that AdvPT differs greatly from AT, which iteratively generates adversarial examples at each iteration of training and continuously updates the target model on the generated adversarial examples, leading to significant computational costs. Conversely, AdvPT fixes the parameters of the image encoder $E$, channeling focus exclusively on updating the learnable vectors at the input of the text encoder $G$. This strategy significantly diminishes the number of learnable parameters. With the image encoder $E$ frozen, the generation of the adversarial examples is only a one-pass process. These examples, once processed through $E$, constitute the adversarial embedding bank $\mathbf{A}$. After this step, the image encoder $E$ is discarded, leaving only the adversarial embedding bank $\mathbf{A}$ for the subsequent prompt tuning.
@@
The design of the adversarial embedding bank presents significant advantages. Primarily, it eliminates the need for redundant forward and backward passes through the image encoder, thereby greatly saving computational time. Moreover, the embedding space’s lower dimensionality compared to the original image space substantially reduces the required computational memory.
@@ 4.2 Learnable Vector Optimization
The next phase in AdvPT involves the construction and optimization of the learnable vectors. Specifically, our method seeks to model textual prompts with learnable vectors $V=[v_{1},v_{2},\ldots,v_{M}]$, optimized by aligning them with adversarial embeddings, thus rectifying the non-robust features of the images utilized by the model. Initially, the text prompts $[v_{1},v_{2},\ldots,v_{M},c_{i}]_{i=1}^{K}$ are fed into the text encoder $G$, producing text embeddings $\mathbf{T}=[w_{1},w_{2},\dots,w_{K}]^{T}\in\mathbb{R}^{L\times K}$. In the fine-tuning phase, each iteration retrieves a mini-batch $\mathbf{B}=[e^{\prime}_{1},e^{\prime}_{2},\dots,e^{\prime}_{b}]\in\mathbb{R}^{b\times L}$ from the adversarial embedding bank $\mathbf{A}$. Subsequently, the similarity scores $\mathbf{S}=\mathbf{B}\mathbf{T}\in\mathbb{R}^{b\times K}$ can be calculated, with each element representing the prediction score in the following manner:
@@
5 Experiments
In this section, we begin by comparing the adversarial robustness of our proposed approach with hand-crafted prompts under both white-box and black-box adversarial attacks. Second, we compare our method with the state-of-the-art input denoising defensive approaches. Additionally, we investigate the trade-off between generalizability and adversarial robustness in the context of prompt tuning. We also discuss the efficiency between our method and AT. Next, we examine the performance of learnable vectors when trained on a specific dataset but evaluated across various distinct datasets. Finally, we carry out an experimental analysis into interpreting the learnable vectors and perform an exhaustive analysis of hyperparameters.
5.1 Experimental Settings
Datasets.
We conduct our study mainly on 8 high-resolution vision datasets: Pets [24], Flowers [22], ImageNet [28], Food101 [1], SUN397 [38], DTD [2], EuroSAT [9], and UCF101 [30]. We adhered to the division of training and testing sets as established in the setup of [47]. For the ImageNet test set, in a manner consistent with prior studies focusing on adversarial attacks [5, 34, 40], we use 1,000 images which are randomly sampled (one image per class). Furthermore, to assess the domain generalization capabilities, we employed four variant datasets of ImageNet, namely ImageNetV2 [27], ImageNet-Sketch [33], ImageNet-A [11], and ImageNet-R [10].
Models.
Adversarial Attacks.
To evaluate adversarial robustness, we introduced both white-box and black-box adversarial attacks. For white-box adversarial attacks, we employed PGD-40 [18], aimed at maximizing KL Divergence in the embedding space, and AutoAttack [3], aimed at maximizing the contrastive loss between image-text pairs, respectively. Regarding black-box attacks, we implemented black-box attack RAP [25].
Adversarial Defenses.
Implementation Details.
Our methodology builds upon the CoOp framework [^1] https://github.com/KaiyangZhou/CoOp. Our training process consists of 5 epochs with a batch size of 512 on ImageNet, and 100 epochs with a batch size of 32 on other datasets. The learnable vectors are optimized via SGD, starting with an initial learning rate of 0.002 for ViT-L/14 and 0.005 for ViT-B/16, and adjusted by cosine annealing. The number of learnable vector $M=32$. To construct the adversarial embedding bank $\mathbf{A}$, we apply the PGD-10 attack with a maximum perturbation of 8/255 over 10 iterations. For white-box adversarial attack on the test set, we utilize PGD-40 with a maximum perturbation of 16/255 over 40 iterations. We conduct black-box adversarial attacks on the test set using RAP for 400 iterations. For the selection of the RAP attack surrogate model, we employ ResNet-50 with torchvision weights for ImageNet, and train an additional fully connected layer on downstream datasets. The hyperparameter $\sigma$ in Super-resolution was set to 0.2. The pre-trained diffusion models in DiffPure is Guided Diffusion [4] and the time step was set to 150.
5.2 Comparison with Vanilla CLIP
Table 1: Accuracy (%) under PGD-40 and RAP attacks: The “( ↑)” indicates the margin by which AdvPT surpasses the vanilla CLIP (hand-crafted prompts).
@@ 5.2 Comparison with Vanilla CLIP
We started our evaluation by comparing AdvPT with the vanilla CLIP model. Using PGD-40 and RAP, we evaluated adversarial robustness in 8 datasets, as indicated in Tab. 1. Our findings reveal that: (1) AdvPT demonstrates improvements over the vanilla CLIP under both PGD-40 and RAP attacks, with the specific improvement quantified in green. (2) While the primary goal of AdvPT is not to enhance generalizability, the empirical finding implies that the enhancement of accuracy emerges as a collateral advantage.
5.3 Comparison with Adversarial Defenses
Table 2: Accuracy (%) under PGD-40 Attack: The “+AdvPT” indicates our method combined with the input denoising defense. The best results are highlighted in bold.
@@ 5.3 Comparison with Adversarial Defenses
Table 3: Accuracy (%) under AutoAttack.
Flowers Pets Food101 SUN397 DTD EuroSAT UCF101 ImageNet
No defense 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
AdvPT 23.3(23.3↑) 7.2(7.2↑) 4.4(4.4↑) 17.6(17.6↑) 28.9(28.9↑) 27.5(27.5↑) 18.5(18.5↑) 11.0(11.0↑)
DiffPure 54.1 78.9 61.1 51.5 35.1 32.9 56.7 55.5
ViT-B/16 +AdvPT 80.3(26.2↑) 84.7(5.8↑) 65.5(4.4↑) 61.4(9.9↑) 60.8(25.7↑) 59.1(26.2↑) 70.8(14.1↑) 60.4(4.9↑)
No defense 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
AdvPT 19.2(19.2↑) 3.3(3.3↑) 3.3(3.3↑) 15.5(15.5↑) 25.7(25.7↑) 29.5(29.5↑) 17.0(17.0↑) 9.0(9.0↑)
DiffPure 63.7 87.5 67.4 51.5 43.4 34.8 64.7 60.7
ViT-L/14 +AdvPT 89.3(25.6↑) 87.9(0.4↑) 72.5(5.1↑) 68.5(17.0↑) 65.4(22.0↑) 63.2(28.4↑) 77.8(13.1↑) 67.9(7.2↑)
As described previously, AdvPT presents an innovative approach to enhance the robustness of image encoders against adversarial attacks by modifying only the textual input. This method is inherently synergistic with visual-modality input denoising defenses. We evaluated its performance against white-box PGD-40, and observed its compatibility with defenses, as delineated in Tab. 2. Significantly, incorporation of AdvPT requires no specialized tuning for the purified images.
Remarkably, the synergy of AdvPT with baseline defense mechanisms sometimes yielded “$1+1>2$” contribution. For example, on the ViT-B/16 model applied to the Flowers dataset, AdvPT alone increases accuracy by 31.0% (from 6.40% to 37.40%), yet when combined with Super-resolution, it further improves the performance of Super-resolution by 46.60% (from 13.80% to 60.4%). In addition to this, we also introduced AutoAttack, which targets the contrastive loss of image-text pairs. The results in Tab. 3 are consistent with those in Tab. 2, indicating that when combined with the state-of-the-art diffusion model-based defense method, DiffPure, our AdvPT achieved enhanced performance. These findings highlight the potential of this innovative synergy strategy to enhance adversarial defense by simultaneously modifying textual and visual inputs, warranting further investigation in future study.
5.4 Generalization-Robustness Trade-off
(a) Flowers
(b) Pets
(c) Food101
(d) SUN397
(e) DTD
(f) EuroSAT
(g) UCF101
(h) ImageNet
Figure 3: AdvPT vs. CoOp on generalization and adversarial robustness.
In this subsection, we discuss the effects of various learning objectives on the learnable vectors within prompt tuning. The primary goal of AdvPT is to enhance the adversarial robustness of the image modality in VLMs. In contrast, we explore whether an objective like CoOp [47], which is fine-tuned on clean images for improved accuracy, affects adversarial robustness differently.
Our comparative analysis of AdvPT and CoOp unveils insights into their generalizability and adversarial robustness, as illustrated in Fig. 3. Our findings are twofold: (1) Intriguingly, AdvPT significantly outperforms CoOp in adversarial robustness, albeit at a slight cost to generalization. This highlights a potential trade-off between adversarial robustness and generalization in prompt tuning, aligning with conclusions drawn from traditional AT [42]. (2) Although AdvPT sacrifices some generalizability, this drawback is mitigated as the model scale increases. Particularly on ViT-L/14, while also enhancing adversarial robustness, the narrowed generalizability gap makes AdvPT highly compatible with the ongoing trend towards scaling up models.
5.5 Comparison with Adversarial Training
Figure 4: Efficiency comparison between AdvPT and AT on Pets.
Figure 5: Efficiency comparison between AdvPT and Fast AT on Pets.
In this subsection, we compare the efficiency of our method, which focuses on fine-tuning only the prompt, against traditional AT. Specifically, we juxtapose AdvPT with PGD-10 AT [18] on the Pets dataset, ensuring that both methods use an equivalent batch size. The comparative results are shown in Fig. 4, including the time taken to compute the adversarial embedding bank $\mathbf{A}$ in the total time reported. We also presented the results of Fast AT [37] in Fig. 5. Although Fast AT is much faster than AT, it still lags significantly behind AdvPT.
Our analysis reveals that AdvPT is more time-efficient than AT, requiring at least an order of magnitude less time. Moreover, it demonstrates a superior enhancement in model performance, outperforming AT by at least one order of magnitude in effectiveness. This efficiency advantage, exceeding 100$\times$ at least, positions AdvPT as a notably superior solution for VLMs.
5.6 Comparison with Linear Prob CLIP
Table 4: Clean accuracy and robust accuracy (PGD-40) of linear prob CLIP.
Flowers Pets Food101 SUN397 DTD EuroSAT UCF101
ViT-B/16 Clean 97.9 91.1 88.4 75.7 77.1 94.3 83.8
PGD 4.8 10.9 5.2 4.8 13.8 9.2 4.1
ViT-L/14 Clean 99.4 94.2 90.9 79.0 80.1 95.9 88.7
PGD 6.8 15.7 12.8 7.7 14.5 21.2 8.0
In this section, we compare AdvPT with linear prob CLIP, which also utilizes additional data, to investigate whether the robustness improvements of AdvPT merely result from additional downstream data, as shown in Tab. 4. By comparing with Tab. 1, while it shows an increase in clean accuracy compared to vanilla CLIP, its robustness is reduced, when compared to AdvPT. This indicates that merely introducing additional downstream data does not directly contribute to enhanced robustness. Furthermore, it also indicates that the enhancements in robustness are not entirely relevant to improvements in accuracy.
5.7 Evaluation on Domain Shift